# Towards Cumulative Abstract Semantics via Handlers

### v1

@@ Abstract.
We consider the problem of modularizing control flow in a generic abstract interpretation framework. A generic abstract interpretation framework is not truly flexible if it does not allow interpreting with different path- and flow-sensitivities, by going forwards or backwards, and over- or under-approximately. Most interpreters inherently intertwine syntax and semantics, making the implementation antagonistic to modularity. Current approaches to modular designs require the use of complex data structures (e.g., monad transformers), providing modularity but often proving unwieldy (e.g., lifts). We observe that leveraging scoped effects within an interpreter facilitates the accumulation of semantic fragments against a fixed syntax. In this paper, we define cumulative abstract semantics, illustrating the potential for creating multiple dynamic evaluators and static analyses from one interpreter. This modularity is achieved by grouping effects into two categories: syntax elimination and domain-semantic introduction handlers. Our contribution shows the benefits of using effects as an instrument for designing a clean, elegant, and modular abstract interpretation framework.
@@ 1. Introduction
Interpreters serve the function of giving semantics to syntax, meaning separating the two is a difficult task. One existing approach is through monads, but the complexity often prevents the full realization of monadic modularity (Brady13). Existing modular monadic approaches use tightly coupled components, traversed with monad transformers, making it difficult to modify elements in isolation (Schrijvers19; Liang95). Monadic encapsulation is represented by layers of stacked monads, each handling specific functionality, relying on monad transformers to couple distinct components. This structure means computations must be manually threaded through adjacent layers.
Recently, a new approach to functional encapsulation has emerged in this space: effects and effect handlers (Plotkin_2013; Brady13; Poulsen23). Several new languages natively support these effects systems, such as Koka (leijen14; leijen17) and Flix (Madsen16). While these languages have explored effect systems for language design and extension, their potential for modularizing abstract interpretation frameworks has remained largely uncharted. At a high level, the distinction between an effect-oriented interface and a monad stack is the ease of composition. However, languages with first-class effects, such as Koka, can infer the effect being triggered and call the corresponding handler automatically based on the inferred effect of the computation (leijen14). Consequently, effect systems allow for the separation of syntax and semantics with effect signatures and effect handlers; to do this, they rely on delimited continuations within handlers, implemented to fill out the type requirements of the effects they handle (Hillerstrom16). By varying the order of applied resumptions in handlers for the same effects, control flow can be transformed (Hillerstrom16; leijen17). As a result the same effectful skeleton or effect interface can express single-path or multiple-path interpretations by changing only one handler.
@@ 2. Cumulative Abstract Semantics
Let us write for the judgment form that says, ”Expression evaluates to integer ,” which is the form a standard, big-step operational semantics. We call an implementation of this judgment form, a substantiated interpreter, as it is a executable function — even if it is effectful:
That is, it may have core language effects, inferred by Koka like divergence div and exceptions exn but otherwise, does not have any unhandled effects.
@@ 2.1. Domain Parametrization and Introduction Handlers
Following abstract interpretation (Cousot77), we first refactor the concrete interpreter from Figure 1 into an abstract interpreter by parametrizing it over an evaluation domain:
@@ 2.2. Interpretation Parametrization and Elimination Handlers
where evaluates and if the result is non-zero, evaluates ; otherwise, it evaluates . In a standard big-step concrete interpreter, the control flow of interpreting will only evaluate one of the branches or as shown in Figure 3. Extending the unsubstantiated interpreter to handle conditionals with an introduction handler
@@ 3. Related Work
To represent effects in Haskell, Kiselyov championed the use of free monads, while also allowing the composition of computations by taking their co-product (Kiselyov15). As a result, evaluation was represented as a fold over the co-product type signature of every effect. This facilitated an extensible effect system, at the cost of manually lifting and injecting each matched effect type. The idea to make an effect based interpreter was theorized and toyed with by Reinders23, but never fully implemented until very recently. In BunkenBurg24 started using effect handlers embedded in Haskell to build a modular interpreter, one of the first of its kind. However, its focus was on concrete interpretation rather than abstract interpretation. This work highlights the potential of embedded effects for modular interpreters, which our approach extends to abstract domains and abstract semantics.
@@
Keidel18 used arrow transformers from category theory to compose abstract domains while preserving soundness in the Sturdy abstract interpretation framework. Their approach provides soundness guarantees for free but is fixed to non-relational domains and requires an implementation of a monolithic, generic interpreter for every language. Ilya13 demonstrated the ability of monads to modularize semantics, allowing them to be combined in unique ways for various interpretations similar to our technique. However this monadic approach carries the same burden of threading computations and context through a dense monadic stack of operations. Michelland24 advanced the usage of monadic modularity with their ITree-based framework for abstract interpretation, composing state and control flow monads for sound meta-theory in Roq. Their approach supports modular analyses (e.g., binding-time analysis) but requires complex transformer stacks. We seek to expand on this idea with our effect based skeleton, which relies on Koka’s handlers for seamless composition single and multi-path analyses, aiming for easier integration.
@@ 4. Conclusion
We presented the groundwork for a modular abstract interpretation framework that leverages algebraic and scoped effects for cumulative semantic definition. Unlike monadic interpreters requiring manual plumbing, row-polymorphic effects automate handler selection, allowing for a seamless sequencing of effectful domain-specific interpretations.

### v2

@@ Abstract.
We consider the problem of modularizing control flow in a generic abstract interpretation framework. A generic abstract interpretation framework is not truly flexible if it does not allow interpreting with different path- and flow-sensitivities, by going forwards or backwards, and over- or under-approximately. Most interpreters inherently intertwine syntax and semantics, making the implementation antagonistic to modularity. Current approaches to modular designs require the use of complex data structures (e.g., monad transformers), providing modularity but often proving unwieldy (e.g., lifts). We observe that leveraging algebraic effects within an interpreter facilitates the accumulation of semantic fragments against a fixed syntax. In this paper, we define cumulative abstract semantics, illustrating the potential for creating multiple dynamic evaluators and static analyses from one interpreter. This modularity is achieved by grouping effects into two categories: syntax elimination and domain-semantic introduction handlers. Our contribution shows the benefits of using effects as an instrument for designing a clean, elegant, and modular abstract interpretation framework.
@@ 1. Introduction
Interpreters serve the function of giving semantics to syntax, meaning separating the two is a difficult task. One existing approach is through monads, but the complexity often prevents the full realization of monadic modularity (Brady, 2013). Existing modular monadic approaches use tightly coupled components, traversed with monad transformers, making it a complex task to modify elements in isolation (Schrijvers et al., 2019; Liang et al., 1995). Monadic encapsulation is represented by layers of stacked monads, each handling specific functionality, relying on monad transformers to couple distinct components. This structure means computations must be manually threaded through adjacent layers.
Recently, a new approach to functional encapsulation has emerged in this space: effects and effect handlers (Plotkin and Pretnar, 2013; Brady, 2013; Bach Poulsen and van der Rest, 2023). Several new languages natively support these effects systems, such as Koka (Leijen, 2014, 2017) and Flix (Madsen et al., 2016). While these languages have explored effect systems for language design and extension, their potential for modularizing abstract interpretation frameworks has remained largely uncharted. At a high level, the distinction between an effect-oriented interface and a monad stack is the ease of composition. However, languages with first-class, polymorphic effects, such as Koka, can infer a triggered effect’s return type and call the corresponding handler automatically (Leijen, 2014). Consequently, effect systems allow for the separation of syntax and semantics with effect signatures and effect handlers; to do this, they rely on delimited continuations within handlers, implemented to fill out the type requirements of the effects they handle (Hillerström and Lindley, 2016). It is important to note that dynamically scoped handlers can cause continuations to capture more than they should when an effect resumes. Grouping effectful operations with multiple resumptions into one effect interface ensures each operations’ handler is in the same scope. As a result, each continuation captures precisely what it should at every resumption. By varying the order of applied resumptions in handlers for the same effects, control flow can be transformed (Hillerström and Lindley, 2016; Leijen, 2017). As a result the same effectful skeleton or effect interface can express single-path or multiple-path interpretations by changing only one handler.
@@ 2. Cumulative Abstract Semantics
Let us write for the judgment form that says, ”Expression evaluates to integer ,” which is the form of a standard, big-step operational semantics. We call an implementation of this judgment form, a substantiated interpreter, as it is a executable function — all effects that must be handled have corresponding handlers, allowing the interpreter to execute programs:
That is, it may have core language effects, inferred by Koka like divergence div and exceptions exn, however these effects do not require handling allowing their presence to preserve our notion of substantiation.
@@ 2.1. Domain Parametrization and Introduction Handlers
Following abstract interpretation (Cousot and Cousot, 1977), we first refactor the concrete interpreter from Figure 1 into an abstract interpreter by parametrizing it over an evaluation domain:
@@ 2.2. Interpretation Parametrization and Elimination Handlers
where evaluates and if the result is non-zero, evaluates ; otherwise, it evaluates . In a standard big-step concrete interpreter, the control flow of interpreting the syntax will only evaluate one of the branches or as shown in Figure 3. Extending the unsubstantiated interpreter to handle conditionals with an introduction handler
@@
This framework also allows changes in control with reuse in the rest of the handlers. For example, a backwards interpretation would only require changing the plus elimination handler to the following:
This new handler evaluates the right hand side before the left hand side, changing the order of evaluation and allowing reuse of the other three handlers.
@@ 3. Related Work
To represent effects in Haskell, Kiselyov championed the use of free monads, while also allowing the composition of computations by taking their co-product (Kiselyov and Ishii, 2015). As a result, evaluation was represented as a fold over the co-product type signature of every effect. This facilitated an extensible effect system, at the cost of manually lifting and injecting each matched effect type. The idea to make an effect based interpreter was theorized and toyed with by Reinders (2023), but never fully implemented until very recently. In Bunkenburg and Wu (2024) started using effect handlers embedded in Haskell to build a modular interpreter, one of the first of its kind. However, its focus was on concrete interpretation rather than abstract interpretation. This work highlights the potential of embedded effects for modular interpreters, which our approach extends to abstract domains and abstract semantics. The Semantic project combined various functional programming components to create a multilingual code differential analysis of GitHub repositories, relying on algebraic effects (Thomson et al., 2022; GitHub, 2025).
@@
Keidel et al. (2018) used arrow transformers from category theory to compose abstract domains while preserving soundness in the Sturdy abstract interpretation framework. Their approach provides soundness guarantees for free but is fixed to non-relational domains and requires an implementation of a monolithic, generic interpreter for every language. Sergey et al. (2013) demonstrated the ability of monads to modularize semantics, allowing them to be combined in unique ways for various interpretations similar to our technique. However this monadic approach carries the same burden of threading computations and context through a dense monadic stack of operations. Michelland et al. (2024) advanced the usage of monadic modularity with their ITree-based framework for abstract interpretation, composing state and control flow monads for sound meta-theory in Roq. Their approach supports modular analyses (e.g., binding-time analysis) but requires complex transformer stacks. We seek to expand on this idea with our effect based skeleton, which relies on Koka’s handlers for seamless composition single and multi-path analyses, aiming for easier integration.
@@ 4. Conclusion
We presented the groundwork for a modular abstract interpretation framework that leverages algebraic effects for cumulative semantic definition. Unlike monadic interpreters requiring manual plumbing, row-polymorphic effects automate handler selection, allowing for a seamless sequencing of effectful domain-specific interpretations.